OpenStudy (anonymous):

Helen works at a local pet shop. She uses 5 kilograms of dog food to feed 3 dogs for 4 days. At the same feeding rate, how many kilograms of dog food will Helen need to feed 12 dogs for 1 week? Steps please!

OpenStudy (stamp):

3 dogs for 4 days = 5 kg 12 dogs for 4 days = 20 kg These rates are for 4 days of care. 4 days * 7/4 = 7 days, so scale up the amount of food for 12 dogs for 4 days. 20 kg * 7/4. 20 kg * 7/4 = 20(7)/4 = 140/4 = 70/2 = 35 kg for 12 dogs for 4 days

OpenStudy (stamp):

Alternate way of breaking down the problem: 3 dogs for 4 days = 5 kg 12 dogs for 4 days = 20 kg 12 dogs for 1 day = 5 kg 12 dogs for 7 days = 35 kg

OpenStudy (stamp):

In my first post, "20 kg * 7/4 = 20(7)/4 = 140/4 = 70/2 = 35 kg for 12 dogs for 4 days" SHOULD INSTEAD READ "20 kg * 7/4 = 20(7)/4 = 140/4 = 70/2 = 35 kg for 12 dogs for 7 days"

OpenStudy (anonymous):

Thank you! :)

OpenStudy (stamp):

Anytime